OpenStudy (anonymous):

differentiate \[q(t)=\frac{ 6t }{ t^2-3t+6 }\]

OpenStudy (anonymous):

So what I am doing is \[\frac{ ((t^2-3t+6)(\frac{ d }{ dx })(6t)) -((6t)(\frac{ d }{ dx })(t^2-3t+6))}{ (t^2-3t+6)^2 }\]

OpenStudy (mathstudent55):